OpenStudy (anonymous):

I need help really fast! please!! A circle has an area of 36π square centimeters. What is the diameter?

Directrix (directrix):

@UmmmHELP The area of a circle is A = pi * r^2 where r is the radius. A = pi * r^2 36 pi = pi * r^2 36 = r^2 So, r = ? Once you get the value of the radius, remember to double to get the diameter.

OpenStudy (anonymous):

Yeah I got it already! thanks(: